26) Calculate the pH of 0.0385 M hypochloric acid [HClO]. [Ka= 3.0*10-8]

explain answer please and show work
Answer #1

                                           HClO      ?            H+ +               ClO-

Initial                                   0.0385                   0                            0

after dissocialtion                 0.0385-x               x                           x

Since Ka is very small so x will also be small.

So, 0.0385-x ?0.0385

Ka = [ClO-][ H+]/[ HClO]

3*10^-8 = x*x/0.0385

x = squareroot(0.0385*3*10^-8)

x = 3.40*10^-5

pH = -log[H+]

      = 4.46